# Southwest Tree: A Low-Memory Data Structure for Partial Accumulations by Non-Commutative Invertible Operations

### v1

@@ 1 Introduction
This paper first outlines the general structure, layout, and properties of the Southwest Tree by defining variables, defining parent/child relationships, defining the values of each node, and providing visualizations of an example. It then describes how to construct the tree and perform the update and accumulate operations while providing Python code samples to do so. Finally, the time and space complexity is analyzed and compared to that of the Segment Tree and Binary Indexed Tree.
@@ 2 Related Work
One data structure in use for this problem is the Segment Tree [1], which works by setting the initial values as leaf nodes of the tree and propagating upwards, combining two adjacent nodes into their parent. This adds nodes to store combined values in addition to the initial values themselves. This does not affect time complexity at large scales, but it does require up to nearly twice as much memory as other options, which could be more important for certain applications or operations. This data structure can be constructed in $O(N)$ time, perform the update and accumulate operations in $O(\log_{2}N)$ time, and requires up to $12N-$ 1 nodes.
The Binary Indexed Tree, or Fenwick Tree [4], is the other leading option for this problem. It works by finding children and parents by simply incrementing or decrementing the last set bit in a node index’s binary representation, while storing the accumulation from the parent index (including) to it’s own index (excluding) of the initial array values. This saves on memory space compared to the Segment Tree. Hence, this data structure can be constructed in $O(N)$ time, perform the update and accumulate operations in $O(\log_{2}N)$ time, and requires $N$ nodes.
@@
Dietz [3], later modified by Ramen, Ramen, and Rao [8], proposes a structure that runs the update and accumulate operations in $\Theta(\log_{2}N/\log_{2}\log_{2}N)$ amortized time. It works by storing “the list at the leaves of a nearly complete tree of branching factor b = $\Theta(\log^{\epsilon}n)$,” where $\epsilon$ is a positive constant less than one. It also uses a two-array scheme and precomputed tables for internal nodes to track children in constant amortized time. However, this scheme requires auxiliary memory greater than a single array of $N$ elements, and each internal node has worst-case time of $O(\log_{2}N)$ to update.
Hon, Sadakane, and Sung [6] review solutions to the similar Searchable Partial Sum problem, which includes the capability of the search opertaion, where $search(j)$ returns the smallest $i$ such that $sum(i)\geq j$. These solutions, while interesting, do not provide improvements on the operations of the Partial Sum problem.
@@ 3 Structure and Layout
The following variables will be defined here and referenced throughout the paper. $I$ is the initial array of $N$ elements of which we want to keep track, which will be modified in-place to represent the Southwest Tree that supports the update and accumulate operations. We will also use $A$ as an array representing the Southwest Tree, modifying $I$ in-place. These two variable represent the same place in memory but are useful distinctions in proofs and explanations. Array indexing begins at 1 in visuals and text descriptions for ease of understanding, but it begins at 0 in code examples for slightly improved memory optimization. We define the $+$ operation on the value types stored in $I$ as a suitable accumulation operation, including non-commutative and invertible, such as matrix multiplication on invertible matrices. We still refer to this operation as a sum throughout the paper, as it makes the use of traditional mathematical symbols, such as $\sum$, available.
$\texttt{update}(i,d)$ adds $d$ to $I[i]$, and $\texttt{accumulate}(i)$ returns the sum of the first $i$ elements in $I$. $i$ will be assumed to be a valid index in the tree, i.e. no less than 1 and no greater than $N$ in one-based indexing. Phantom nodes are nodes that do not exist in memory but whose indices are used for tree traversal. These are described in more detail in Sec. 3.3.
This paper refers to an ongoing example where $I$ initially consists of the odd numbers from 1 to 17, inclusive, as shown in Fig. 1. These elements will be theoretically arranged in the Southwest Tree as shown in Fig. 2(a), but are actually stored by modifying $I$ to reflect the tree, as shown in Fig. 2(b).
@@ 3.2 Node Values
Since the tree is built bottom-up (see Sec. 3.3), the child elements stored in $A$ will already have the values of the tree.
@@ 3.3 Building the Tree
The algorithm begins at the root index, which will be $12^{\lfloor\log_{2}N\rfloor+1}-$ 1. It then traverses until it reaches a leaf node index, which will be when the node height is 1. The node value of the leaf nodes will be the initial element at the node’s index. The algorithm then traverses back up the tree, where each node value will be the sum of its children plus the initial element at the node’s index. Note that the following code examples build upon each other, so assume that each code example has access to the code listed in above examples.
Although the algorithm will traverse through indices up to the root index, it will only store values in the original array and therefore does not store values at extra indices. We call the indices that are traversed but greater than the size of $I$ phantom indices, and we call nodes at these indices phantom nodes. That is, they serve to help traversal, but only in mathematical operations.
@@
Given an initial array of values, $I$, modified according to Sec. 3.3, $A[i]=\sum_{j=i-2^{h_{i}}+2}^{i}I[j]$, where $h_{i}$ is the height of the node at index $i$.
@@
We begin to see the pattern, then, stemming from the fact that all nodes contain the accumulation of values from its left and right subtrees, along with its own index. Both subtrees accumulate values from all indices beginning with the leftmost leaf up to its own index. Since the leftmost leaf of the right subtree has one greater index than the root of the left subtree, all nodes accumulate all indices from the leftmost leaf up to its own index. In general, for any node at index $i$ with height $h_{i}$,
@@ 4 Operations
The following descriptions use the variable $curr$ to track the index of the current node in the recursive traversal. If the desired index $i$ is less than or equal to the index of the left child, then traverse into the left subtree. Otherwise, traverse into the right subtree. Recurse until you reach the node where $curr==i$. The helper function in Listing LABEL:lst:getchild finds the index of the subtree into which the recursion should traverse, as just described.
Listing 2: A Python example of a helper function that returns the index of the left or right subtree into which the recursion should traverse. $i$ is the desired index  $curr$ is the current index  and $left$ is the difference in index to the left child.
@@ 4.1 Accumulate
The $\texttt{accumulate}(i)$ operation, shown in Listing LABEL:lst:accumulate, is performed by traversing the tree beginning at the root node. If $i$ is equal to $curr$, accumulate the value of $A[curr]$ to the total and break out of the traversal. If $i$ is less than or equal to the index of the left child, then traverse into the left subtree. Otherwise, accumulate the value of the left child to the total and traverse into the right subtree. There is no need to check for phantom nodes here, because if the left child is phantom, then $i$ must be less than its index.
@@
The $\texttt{accumulate}(i,delta)$ of the Southwest Tree add nothing when traversing into the left subtree, while traversals into the right subtree add the value of the sibling of the subtree’s root. Therefore, let us consider subarrays of $T$ consisting of consecutive $1$s followed by a single 00. Let the traversal step of traversals into right subtrees be $R=[r_{1},r_{2},\dots,r_{\ell}]=[a:1\leq a\leq k\wedge t_{a}=0]$, where $r_{b}$ indicates the $b$th element of $R$ and $\ell$ is the final index of $R$.
Let $\mathcal{P}(i)$ be the parent of the node at index $i$, with $\mathcal{P}_{r}(i)$ being the parent of a right child. By Eq. (2), given index $i$ of a right child,
@@
The first $t_{a}=0$ will be preceded by a sequence of zero or more $1$s, traversing left from the root node until reaching the height $h_{r}-r_{1}$. Hence, the first operand, detailed in Appendix A, will be
@@
The $m$th $t_{a}=0$ will traverse left, beginning at $\mathcal{I}(r_{m-1})$, until reaching the height $h_{r}-r_{m}$. Hence, the $m$th operand, detailed in Appendix B, will be
@@ 4.2 Update
Given an initial array of values, $I$, modified to an array $A$ according to Sec. 3.3, the function $\texttt{update}(i,\delta)$ described in Sec. 4.2 correctly sets $A[i]=A[i]+\delta$ while maintaining the order of operands.
@@
Say that $inv(A[c])+A[c]$ for any index $c$ yields the identity of the given operator $+$. Each step of the traversal begins by inverting the values of the root node’s children, leaving only $I[c]$. In order to maintain the order of non-commutative operands, we add the inverse of the children in reverse order from Eq. (4), i.e., we set $A[c]=inv(A[c-1])+inv(A[c-2^{h_{c}-1}])+A[c]$.
$A[c]$ $=A[c]=inv(A[c-1])+inv(A[c-2^{h_{c}-1}])+A[c-2^{h_{c}-1}]+A[c-1]+I[c]$ (16)
@@ 5 Complexity Comparison
This section describes the time and space complexity for the Southwest tree and highlights how it outperforms the Segment Tree, which is generally considered the leading solution for non-commutative partial accumulations. The Southwest Tree performs with the same time complexity while reducing the space requirements, so long as the accumulation operation is invertible.
@@ 5.1.2 Build
To build the tree, one simply needs to traverse each node twice, and the tree will have $12^{\lfloor\log_{2}N\rfloor+1}-$ 1 nodes. In the best case, $N=2^{x}-1$ for any integer $x\in\mathbb{N}$, where $\mathbb{N}$ is the set of natural numbers. In this case, the tree will have $2^{\lfloor\log_{2}(2^{x}-1)\rfloor+1}-1=2^{x-1+1}-1=2^{x}-1=N$, having zero phantom nodes.
In the worst case, $N=2^{x}$ for any integer $x\in\mathbb{N}$. In this case, the tree will have $12^{\lfloor\log_{2}(2^{x})\rfloor+1}-1=2^{x+1}-1=2N+$ 1, having $12N+1-N=N-$ 1 phantom nodes.
@@ 5.2 Space
Since both the Southwest Tree and Segment Tree recursively traverse the tree to the bottom in the worst case, they equally require $O(\log_{2}N)$ stack space during operations.
However, although both of these trees have $O(N)$ space complexity to store the values, the Southwest Tree modifies the input array in-place and does not require any additional space for values. This beats the Segment Tree, which needs $O(N)$ additional space to store values on top of the initial array.
@@ 6 Conclusion
The first $t_{a}=0$ will be preceded by a sequence of zero or more $1$s, traversing left from the root node until reaching the height $h_{r}-r_{1}$. Hence, the first operand, by Lemma 3.1, will be
@@
$=\sum_{j=2^{h_{r}-r_{1}+1}-1-2^{h_{r}-r_{1}+1}+2}^{\mathcal{S}(\mathcal{I}(r_{1}))}I[j]$
$=\sum_{j=1}^{\mathcal{S}(\mathcal{I}(r_{1}))}I[j],$
@@
$2^{h_{r}}-1-\sum_{j=1}^{r_{1}-1}2^{h_{r}-j}$ $=2^{h_{r}-r_{1}+1}-1$ (20)
because $12^{h_{r}}-$ 1 can be thought of as a bit string of length $h_{r}$ of all $1$s, and subtracting $\sum_{j=1}^{r_{1}-1}2^{h_{r}-j}$ removes $r_{1}-1-j+1=r_{1}-1$ of the most significant bits, resulting in a bit string of length $h_{r}-(r_{1}-1)=h_{r}-r_{1}+1$ of all $1$s.
@@
The $m$th $t_{a}=0$ will traverse left, beginning at $\mathcal{I}(r_{m-1})$, until reaching the height $h_{r}-r_{m}$. Hence, the $m$th operand, by Lemma 3.1, will be
@@
$=\sum_{j=\mathcal{I}(r_{m})+1-2^{h_{(\mathcal{I}(r_{m}))}-1}-2^{h_{r}-r_{m}}+2}^{\mathcal{S}(\mathcal{I}(r_{m}))}I[j]$
@@
$=\sum_{j=\mathcal{I}(r_{m-1})-\left(\sum_{j=r_{m-1}+1}^{r_{m}-2}2^{h_{r}-j}\right)-2^{h_{r}-r_{m}+2}+2}^{\mathcal{S}(\mathcal{I}(r_{m}))}I[j]$
$=\sum_{j=\mathcal{I}(r_{m-1})-\left(\sum_{j=r_{m-1}+1}^{r_{m}-3}2^{h_{r}-j}\right)-2^{h_{r}-r_{m}+3}+2}^{\mathcal{S}(\mathcal{I}(r_{m}))}I[j]$
=$\dots$= …
@@
If $t_{k}=0$, then $\ell=k$ and $\mathcal{I}(k)=\mathcal{S}(\mathcal{I}(r_{\ell}))+2^{h_{r}-r_{\ell}}-1$ and

### v2

@@ 1 Introduction
This paper first outlines the general structure, layout, and properties of the Southwest Tree by defining variables, defining parent/child relationships, defining the values of each node, and providing visualizations of an example. It then describes how to construct the tree and perform the update and accumulate operations while providing Python code samples to do so. Finally, the time and space complexity is analyzed and compared to that of the Segment Tree.
@@ 2 Related Work
One data structure in use for this problem is the Segment Tree, which works by setting the initial values as leaf nodes of the tree and propagating upwards, combining two adjacent nodes into their parent [1]. This adds nodes to store combined values in addition to the initial values themselves. This does not affect time complexity at large scales, but it does require up to nearly twice as much memory as other options, which could be more of a detriment for certain memory-limited applications or operations. This data structure can be constructed in $O(N)$ time, perform the update and accumulate operations in $O(\log_{2}N)$ time, and requires up to $12N-$ 1 nodes.
The Binary Indexed Tree, or Fenwick Tree, is the other leading option for this problem [4]. It works by finding children and parents by simply incrementing or decrementing the last set bit in a node index’s binary representation, while storing the accumulation from (including) the parent index to (excluding) it’s own index of the initial array values. This saves on memory space compared to the Segment Tree. Hence, this data structure can be constructed in $O(N)$ time, perform the update and accumulate operations in $O(\log_{2}N)$ time, and requires $N$ nodes.
@@
Dietz, later modified by Ramen, Ramen, and Rao, proposes a structure that runs the update and accumulate operations in $\Theta(\log_{2}N/\log_{2}\log_{2}N)$ amortized time [3, 8]. It works by storing “the list at the leaves of a nearly complete tree of branching factor b = $\Theta(\log^{\epsilon}n)$,” where $\epsilon$ is a positive constant less than one. It also uses a two-array scheme and precomputed tables for internal nodes to track children in constant amortized time. However, this scheme requires auxiliary memory greater than a single array of $N$ elements, and each internal node has worst-case time of $O(\log_{2}N)$ to update.
Hon, Sadakane, and Sung review solutions to the similar Searchable Partial Sum problem, which includes the capability of the search opertaion, where $search(j)$ returns the smallest $i$ such that $sum(i)\geq j$ [6]. These solutions, while interesting, do not provide improvements on the operations of the Partial Sum problem.
@@ 3 Structure and Layout
The following variables will be defined here and referenced throughout the paper. $I$ is the initial array of $N$ elements of which we want to keep track, which will be modified in-place to represent the Southwest Tree that supports the update and accumulate operations. We will also use $A$ as an variable name representing the in-place modified values of $I$ according to the construction of the Southwest Tree. These two variables represent the same memory address but are useful distinctions in proofs and explanations. Array indexing begins at 1 in visuals and text descriptions for ease of understanding, but it begins at 0 in code examples for slightly improved memory optimization. We define the $+$ operation on the value types stored in $I$ as a suitable accumulation operation, including non-commutative and invertible, such as matrix multiplication on invertible matrices. We still refer to this operation as a sum throughout the paper, as it makes the use of traditional mathematical symbols, such as $\sum$, available.
$\texttt{update}(i,d)$ adds $d$ to $I[i]$, and $\texttt{accumulate}(i)$ returns the sum of the first $i$ elements in $I$. $i$ will be assumed to be a valid index in the tree, i.e., no less than 1 and no greater than $N$ in one-based indexing. Phantom nodes are nodes that do not exist in memory but whose indices are used for tree traversal. These are described in more detail in Sec. 3.3.
This paper refers to an ongoing example where $I$ initially consists of the odd numbers from 1 to 17, inclusive, as shown in Fig. 1. These elements will be theoretically arranged in the Southwest Tree as shown in Fig. 2(a), but are actually stored by modifying $I$ to reflect the tree, as shown in Fig. 2(b). Note that the code examples throughout this paper build upon each other, so assume that each code example has access to the code listed in previous examples.
@@ 3.2 Node Values
Since the tree is built bottom-up (see Sec. 3.3), the child elements of $A[i]$ will already have the modified values of the tree.
@@ 3.3 Building the Tree
The algorithm begins at the root index, which will be $12^{\lfloor\log_{2}N\rfloor+1}-$ 1. It then executes a depth-first traversal until it reaches a leaf node index, which will be when the node height is 1. The node value of the leaf nodes will be the initial element at the node’s index. The algorithm then traverses back up the tree, where each node value will be the sum of its children plus the initial element at the node’s index.
Although the algorithm will traverse through indices up to the root index, it will only store values in the original array and therefore does not store values at extra indices. We call the indices that are traversed but greater than $N$ phantom indices, and we call nodes at these indices phantom nodes. That is, they serve to help traversal, but only in mathematical operations.
@@
Given an initial array of values, $I$, modified to an array $A$ according to Sec. 3.3, the value at index $i$ will be $A[i]=\sum_{j=i-2^{h_{i}}+2}^{i}I[j]$ while maintaining the order of operands, where $h_{i}$ is the height of the node at index $i$.
@@
We begin to see the pattern, then, stemming from the fact that all nodes contain the accumulation of values from its left and right subtrees, along with its own index. Both subtrees accumulate values from all indices beginning with the leftmost leaf up to its own index. In general, for any node at index $i$ with height $h_{i}$,
@@
Note here that the leftmost leaf of the right subtree has one greater index than the root of the left subtree.
@@ 4 Operations
The following descriptions use the variable $curr$ to track the index of the current node in the recursive traversal. If the desired index $i$ is less than or equal to the index of the left child, then traverse into the left subtree. Otherwise, traverse into the right subtree. Recurse until you reach the node where $curr==i$. The helper function in Listing LABEL:lst:getchild finds the root index of the subtree into which the recursion should traverse.
Listing 2: A Python example of a helper function that returns the root index of the left or right subtree into which the recursion should traverse. $i$ is the desired index and $left$ is the difference in index to the left child.
@@ 4.1 Accumulate
The $\texttt{accumulate}(i)$ operation, shown in Listing LABEL:lst:accumulate, is performed by traversing the tree beginning at the root node. If $i$ is equal to $curr$, accumulate the value of $A[curr]$ to the total and break out of the traversal. If $i$ is less than or equal to the index of the left child, then traverse into the left subtree. Otherwise, accumulate the value of the left child to the total and traverse into the right subtree. There is no need to check for phantom nodes here, because if the left child is phantom, then $i$ must be less than its index, hence traversing into the tree with no attempt to accumulate a phantom value.
@@
The $\texttt{accumulate}(i,\delta)$ of the Southwest Tree adds nothing when traversing into the left subtree, while traversals into the right subtree add the value of the root of the left subtree. Therefore, let us consider subarrays of $T$ consisting of consecutive $1$s followed by a single 00, which dictate each operand from a rightward traversal. Let the traversal step of traversals into right subtrees be $R=[r_{1},r_{2},\dots,r_{\ell}]=[a:1\leq a\leq k\wedge t_{a}=0]$, where $r_{b}$ indicates the $b$th element of $R$ and $\ell$ is the final index of $R$.
Let $\mathcal{P}(i)$ be the parent of the node at index $i$, with $\mathcal{P}_{r}(i)$ being the parent of a right child, i.e., the left child of its parent. By Eq. (2), given index $i$ of a right child,
@@
The first $t_{a}=0$ will be preceded by a sequence of zero or more $1$s, traversing left from the root node. Hence, the first operand, detailed in Appendix A, will be
@@
The $m$th $t_{a}=0$ will traverse left from the node at index $\mathcal{I}(r_{m-1})$. Hence, the $m$th operand, detailed in Appendix B, will be
@@ 4.2 Update
Given an initial array of values, $I$, modified to an array $A$ according to Sec. 3.3, the function $\texttt{update}(i,\delta)$ described in Sec. 4.2 correctly sets $I[i]=I[i]+\delta$ while maintaining the order of operands.
@@
Say that $inv(A[c])+A[c]$ for any index $c$ yields the identity of the given operator $+$. Each step of the traversal begins by inverting the values of the root node’s children, leaving only $I[curr]$. In order to maintain the order of non-commutative operands, we add the inverse of the children in reverse order from Eq. (4), i.e., we set $A[c]=inv(A[c-1])+inv(A[c-2^{h_{c}-1}])+A[c]$.
$A[c]$ $=inv(A[c-1])+inv(A[c-2^{h_{c}-1}])+A[c-2^{h_{c}-1}]+A[c-1]+I[c]$ (16)
@@ 5 Complexity Comparison
This section describes the time and space complexity for the Southwest Tree and highlights how it outperforms the Segment Tree, which is generally considered the leading solution for non-commutative partial accumulations. The Southwest Tree performs with the same time complexity while reducing the space requirements.
@@ 5.1.2 Build
To build the tree, one simply needs to traverse each node twice, and the tree will have $12^{\lfloor\log_{2}N\rfloor+1}-$ 1 nodes. The best case is when $N=2^{x}-1$ for any integer $x\in\mathbb{N}$, where $\mathbb{N}$ is the set of natural numbers. In this case, the tree will have $2^{\lfloor\log_{2}(2^{x}-1)\rfloor+1}-1=2^{x-1+1}-1=2^{x}-1=N$, having zero phantom nodes. In the worst case, $N=2^{x}$ for any integer $x\in\mathbb{N}$. In this case, the tree will have $12^{\lfloor\log_{2}(2^{x})\rfloor+1}-1=2^{x+1}-1=2N-$ 1, having $12N-1-N=N-$ 1 phantom nodes.
@@ 5.2 Space
Since both the Southwest Tree and Segment Tree recursively traverse the tree to the bottom in the worst case, they equally require $O(\log_{2}N)$ stack space during operations. However, although both of these trees have $O(N)$ space complexity to store the values, the Southwest Tree modifies the input array in-place and does not require any additional space for values. This beats the Segment Tree, which needs $O(N)$ additional space to store values on top of the initial array.
@@ 6 Conclusion
The first $t_{a}=0$ will be preceded by a sequence of zero or more $1$s, traversing left from the root node. Hence, the first operand, by Lemma 3.1, will be
@@
$=\sum_{j=2^{h_{r}}-1-2^{h_{r}-r_{1}+1+(r_{1}-1-1+1)}+2}^{\mathcal{S}(\mathcal{I}(r_{1}))}I[j]$
$=\sum_{j=2^{h_{r}}-1-2^{h_{r}}+2}^{\mathcal{S}(\mathcal{I}(r_{1}))}I[j]$
$=\sum_{j=1}^{\mathcal{S}(\mathcal{I}(r_{1}))}I[j]$
@@
$\left(\sum_{j=1}^{r_{1}-1}2^{h_{r}-j}\right)+2^{h_{r}-r_{1}+1}$ $=2^{h_{r}-r_{1}+1+(r_{1}-1-1+1)}$ (20)
because the summation $\sum_{j=1}^{r_{1}-1}2^{h_{r}-j}$ adds the number of operands of the summation, $(r_{1}-1)-(1)+1$, to the exponent of $2^{h_{r}-r_{1}+1}$, as one can see by combining the operands one by one as follows:
$\left(\sum_{j=1}^{r_{1}-1}2^{h_{r}-j}\right)+2^{h_{r}-r_{1}+1}$ $=\left(\sum_{j=1}^{r_{1}-2}2^{h_{r}-j}\right)+2^{h_{r}-r_{1}+2}$ (21)
$=\left(\sum_{j=1}^{r_{1}-3}2^{h_{r}-j}\right)+2^{h_{r}-r_{1}+3}$
=$\dots$= …
@@
The $m$th $t_{a}=0$ will traverse left from the node at index $\mathcal{I}(r_{m-1})$. Hence, the $m$th operand, following similar logic to Sec. A, will be
@@
$=\sum_{j=\mathcal{I}(r_{m})+1-2^{h_{\mathcal{P}_{r}(\mathcal{I}(r_{m}))}-1}-2^{h_{r}-r_{m}}+2}^{\mathcal{S}(\mathcal{I}(r_{m}))}I[j]$
@@
If $t_{k}=0$, then $r_{\ell}=k$ and $\mathcal{I}(k)=\mathcal{S}(\mathcal{I}(r_{\ell}))+2^{h_{r}-r_{\ell}}-1$ and